Solve the system by addition.
5x – 3y = 13
4x – 3y = 11

Answer by 303795(562)

5x – 3y = 13
4x – 3y = 11
Subtract the second equation from the first to get 1x = 2. Substitute x=2 into the first equation to get
5 x 2 - 3y = 13
10 - 3y = 13 Subtract 10 from each side
10 - 3y - 10 = 13 - 10
-3y = 3
So y = -1
x = 2 and y = -1